Q.4 A random vector (X, Y, Z) has joint density given by

. f (x, y,z) = k exp [− 1/2 (2x2 − 2xy + y2 + z2 + 2x − 6y) .



1. Compute k.

2. Compute the expectations P(X), P(Y ) and P(Z).

3. Compute the density of the random vector (X, Z).

4. Compute the correlation coeffificient between X and Z and between X and Y .

5. Let W = X + Z; compute the probability density of W.
1
Expert's answer
2021-07-16T09:16:04-0400

"f (x, y, z) = k \exp [− (2x^2 − 2xy + y^2 + z^2 + 2x − 6y)/2]"


"f (x, y, z) = k \exp [− ((x-2)^2 − (y-x-3)^2 + z^2 -13)/2]"


"f (x, y, z) = k \exp [− (2(x-(y-1)/2)^2 + (y-5)^2/2 + z^2 -13)/2]"


(1) "1=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}f(x,y,z)dxdydz"


"=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}k \exp [− ((x-2)^2 − (y-x-3)^2 + z^2 -13)/2]dxdydz"


"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}\exp(- z^2/2)dz\int\limits_{-\infty}^{+\infty}\exp (− (x-2)^2/2)\left(\int\limits_{-\infty}^{+\infty}\exp(−(y-x-3)^2 /2)dy\right)dx"

"=k(2\pi)^{3/2}e^{-13/2}"

Therefore, "k=(2\pi)^{-3/2}e^{13/2}".


(2) "E(X)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}xf(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}\exp(- z^2/2)dz\int\limits_{-\infty}^{+\infty}x\exp (− (x-2)^2/2)\left(\int\limits_{-\infty}^{+\infty}\exp(−(y-x-3)^2 /2)dy\right)dx="

"=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}x\exp (− (x-2)^2/2)dx=2"


"E(Y)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}yf(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}\exp(- z^2/2)dz\int\limits_{-\infty}^{+\infty}\exp (− (x-2)^2/2)\left(\int\limits_{-\infty}^{+\infty}y\exp(−(y-x-3)^2 /2)dy\right)dx="

"=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}(x+3)\exp (− (x-2)^2/2)dx=5"


"E(Z)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}zf(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}z\exp(- z^2/2)dz\int\limits_{-\infty}^{+\infty}\exp (−(x-2)^2/2)\left(\int\limits_{-\infty}^{+\infty}\exp(−(y-x-3)^2 /2)dy\right)dx="

"=0"


(3)

"f_{X,Z} (x, z)=\int\limits_{-\infty}^{+\infty}f(x,y,z)dy=\int\limits_{-\infty}^{+\infty} k \exp [− ((x-2)^2 − (y-x-3)^2 + z^2 -13)/2]dy="

"=\frac{x+3}{2\pi}\exp [− ((x-2)^2 + z^2 )/2]"


"V(X)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}(x-2)^2f(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}e^{- z^2/2}dz\int\limits_{-\infty}^{+\infty}(x-2)^2 e^{−(x-2)^2/2} \left(\int\limits_{-\infty}^{+\infty}e^{−(y-x-3)^2 /2}dy\right)dx=1"


"V(Y)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}(y-5)^2f(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}e^{- z^2/2}dz\int\limits_{-\infty}^{+\infty} e^{−(x-2)^2/2} \left(\int\limits_{-\infty}^{+\infty}(y-5)^2e^{−(y-x-3)^2 /2}dy\right)dx="

"=\frac{1}{2\pi}\int\limits_{-\infty}^{+\infty} e^{−(x-2)^2/2} \left(\int\limits_{-\infty}^{+\infty}(y+x-2)^2e^{−y^2 /2}dy\right)dx="

"=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty} (1+(x-2)^2)e^{−(x-2)^2/2}dx="

"=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty} (1+x^2)e^{−x^2/2}dx=2"


(4) "Cov(X, Z)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}z(x-2)f(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}ze^{- z^2/2}dz\int\limits_{-\infty}^{+\infty}(x-2) e^{−(x-2)^2/2} \left(\int\limits_{-\infty}^{+\infty}e^{−(y-x-3)^2 /2}dy\right)dx=0"


Therefore, "Corr(X,Z)=Cov(X,Z)/(\sigma_X\sigma_Z)=0"


"Cov(X, Y)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}(x-2)(y-5)f(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}e^{- z^2/2}dz\int\limits_{-\infty}^{+\infty}(x-2) e^{−(x-2)^2/2} \left(\int\limits_{-\infty}^{+\infty}(y-5)e^{−(y-x-3)^2 /2}dy\right)dx="

"=\frac{1}{2\pi}\int\limits_{-\infty}^{+\infty}(x-2) e^{−(x-2)^2/2} \left(\int\limits_{-\infty}^{+\infty}(y+x-2)e^{−y^2 /2}dy\right)dx="

"=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}(x-2)^2 e^{−(x-2)^2/2}dx=1"


"\sigma_X=\sqrt{V(X)}=1"

"\sigma_Y=\sqrt{V(Y)}=\sqrt{2}"


"Cov(X, Y)=Corr(X,Y)/(\sigma_X\sigma_Y)=1/(1\cdot\sqrt{2})=\sqrt{2}/2"


5."W=X+Z"

"P(W<w)=\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{+\infty}I_{x+z<w}(x,y,z)f(x,y,z)dxdydz="

"=ke^{-13/2}\int\limits_{-\infty}^{+\infty}e^{- z^2/2}\int\limits_{-\infty}^{w-z}e^{−(x-2)^2/2} \int\limits_{-\infty}^{+\infty}e^{−(y-x-3)^2 /2}dydxdz="

"=\frac{1}{2\pi}\int\limits_{-\infty}^{+\infty}e^{- z^2/2}\int\limits_{-\infty}^{w-z}e^{−(x-2)^2/2} dxdz"


"f_W(w)=\frac{d}{dw}P(W<w)=\frac{1}{2\pi}\frac{d}{dw}\int\limits_{-\infty}^{+\infty}e^{- z^2/2}\int\limits_{-\infty}^{w-z}e^{−(x-2)^2/2} dxdz="

"=\frac{1}{2\pi}\int\limits_{-\infty}^{+\infty}e^{- z^2/2}e^{−(w-z-2)^2/2}dz="

"=\frac{1}{2\pi}\int\limits_{-\infty}^{+\infty}e^{-(2z^2-2zw+w^2+4z-4w+4)/2}dz="

"=\frac{1}{2\pi}\int\limits_{-\infty}^{+\infty}e^{-(z-w/2+1)^2}e^{-(w-2)^2/4}dz="

"=\frac{1}{2\pi}e^{-(w-2)^2/4}\int\limits_{-\infty}^{+\infty}e^{-z^2}dz=\frac{1}{2\sqrt{\pi}}e^{-(w-2)^2/4}"